OpenStudy (anonymous):

The 5th term in a geometric sequence is 160. The seventh term is 40. What are the possible values for the sixth term? I think it's 80? Half of 160 is 80, and half of 80 is 40..?

OpenStudy (anonymous):

The sequence is decreasing. Let sixth term be X. So you Have 160, X, 40 Introduce another variable k as the common multiple. So X=k(160) and 40=k(X) Single out k for both equations and equate

OpenStudy (anonymous):

80's right

OpenStudy (anonymous):

Can you help me with more?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Great, it'll take me a moment to find one I need help with.

OpenStudy (anonymous):

What is the equation of a parabola with the vertex at the origin and focus (-5,0)

OpenStudy (anonymous):

@ArsalanAijaz

OpenStudy (anonymous):

parabola or hyperbola?

OpenStudy (anonymous):

Parabola

OpenStudy (anonymous):

It should be something like this.. But I dont know :( \[y=ax ^{2}+bx\]

OpenStudy (anonymous):

That's okay. can you write an equation for the translation of y=6/x that has the asymptotes x=4 and y=5?

OpenStudy (anonymous):

\[y=6\div(x-4)\] I'm forgetting what to do with the horizontal asymptote :(

OpenStudy (anonymous):

I think it's \[y=\frac{ 6 }{ x-4 }+5\]

OpenStudy (triciaal):

regarding the original question the choices means more than one the answer would be 80 and negative 80 the common ratio can be 1/2 or -1/2 the nth term in a geometric progression is given by a1*r^n-1 where a1 is the first term simplified we get a7/a6 = a6/a5 a7*a5 = a6^2 when you take the square root you always have a positive and a negative

OpenStudy (triciaal):